# Joint Attention Mechanism Learning to Facilitate Opto-physiological Monitoring during Physical Activity

### v1

@@ 2.2.2 Cross-attention Mechanism for Multi-sensor Fusion
A self-attention mechanism was proposed as a core component of the transformer model [12]. The self-attention accepts a tensor comprising sequential data and establishes its self-correlation. Specifically, it involves linear projections of the input $X$ into queries ($Q$), keys ($K$), and values ($V$) using different learned weight matrices for each attention head $i$:
$Q(i)=W{{}_{q}}(i)X,K(i)=W{{}_{k}}(i)X,V(i)=W{{}_{v}}(i)X$ (2)
where here $1\leq i\leq[h=8]$ and $X$ is a concatenation of the encoded network inputs (i.e. the encoder output). The scaled dot-product attention ($A$) is then applied to these projections as
$A(i)=softmax((Q(i)K(i)^{T})/\sqrt{d_{k}})V(i)$ (3)
where $d{{}_{k}}$ is the dimensionality of the queries and keys, and the division by $\sqrt{d_{k}}$ is done to prevent the dot product from growing too large as the dimensionality increases.
In contrast to the self-attention mechanism that correlates the input $X$ with itself, the multi-head cross-attention mechanism is proposed for multi-sensor fusion. They are two inputs $X_{1}$ and $X_{2}$, where $X_{1}$ represents query $Q$ tensors and $X_{2}$ represents the key $K$ and value $V$ tensors as shown in (2) and (3). In the proposed attention mechanism, these tensors are related to different sensor inputs, including encoded noisy PPG and encoded triaxial ACC with related VEL feature maps after four downsample blocks of the generator. The encoded feature maps are fed into the attention mechanism, as shown in Fig. 2. Specifically, the proposed attention mechanism takes the generated PPG encoded feature map processed by a (feature pooling) $1\times 1$ convolution layer to create the query $Q$. Also, the key $K$ and value $V$ pair is formed through a similar $1\times 1$ convolution layer of generated combined encoded feature map with triaxial ACC and the related VEL. The $K$ and $V$ vectors are used to calculate the attention weights, indicating to the model which parts of the MAs to focus on. The $Q$ vector is the information that these attention weights will be applied to, with the aim of highlighting the MA features in the PPG encoded feature map that are most relevant to the ACC and VEL combined encoded feature map.
@@ 2.4 Loss function
The discriminator binary cross entropy loss function is [16]
@@ 3 Selective Measurement Protocols
Three datasets provided by Loughborough University (in-house LU dataset), the IEEE Signal Processing Cup 2015 (IEEE-SPC dataset) and the publicly available PPG dataset (PPGDalia dataset) with their ethical approvals are experimented with the proposed AM-GAN.
@@
The IEEE-SPC dataset features a two-channel PPG signal using green (515 nm) LEDs, a triaxial ACC signal, and an ECG reference, collected from 12 male subjects aged between 18 and 35 years. These datasets were sampled at a frequency of 125 Hz [4]. The IEEE-SPC includes the training and testing datasets. In the training datasets, 12 subjects ran at varying speeds. Furthermore, 10 subjects (aged 19-58 years) performed intensive arm movements, i.e., boxing, in the testing datasets.
The PPGDalia includes synchronised recordings of ECG, wrist-worn PPG, and acceleration data from 15 subjects. Each subject contributed approximately two hours of recorded data. A consistent time shift between PPG and triaxial ACC signals was identified and manually fixed [17].
@@ 4.1 Experimental Setup
An eight-second sliding window technique is applied to analyse both the input and reference PPG signals, advancing one second at each step. All input and reference PPG signals are downsampled to 32 $Hz$. During the experiments, the proposed AM-GAN undergoes training for 100 epochs, with both the generator and discriminator networks being optimised concurrently using the Adam optimiser at a learning rate of 0.0002. The input PPG data is pre-processed with an eighth-order Butterworth band-pass filter with band edges set at 0.2 $Hz$ and 6.5 $Hz$. In all experiments, HR and RR are computed using the extraction algorithms defined in [15].
@@ 4.2 Experimental Results
[…]
To validate the error of the PPG signal generated by the AM-GAN method with respect to the reference PPG signal, the Pearson correlation ($R$) was applied for quantifying waveform, as shown in (9), where $G(p(t),a_{x,y,z}(t))$ is the zero-mean generated PPG signal and $s_{ref}(t)=MR(p(t),a_{x,y,z}(t))$ is the zero-mean reference signal, for each vector $p(t)$ in the test set, $\mathcal{P}_{Te}$. An eight-second sliding time window was employed for the calculation of the waveform quality. The waveform quality index ($R$), for the treadmill exercise test cases of subject F01, is shown in Fig. 5(a). The mean Pearson correlation ($\overline{\langle R\rangle}$) for all LU testing subjects is illustrated in Fig. 5(b). It is clear the AM-GAN outputs are highly correlated with the reference signals with a high level of consistency across the entire set of test vectors and for all testing subjects, and the $Mean$ value of all testing subjects’ waveform quality ($\overline{\langle R\rangle}$) is 0.9522.
@@
The outcomes of continuous HR monitoring during treadmill and cycling exercises are shown in Fig. 6(a) and 6(b), where they are compared to the reference values from a commercial device. In Fig. 6, Polar refers to the reference heart rate recorded with a Polar Bluetooth Smart chest strap. In both figures, the MA-removal reference uses our previously published algorithm [15]. The accuracy in capturing HR is consistent and closely matches the provided reference.
@@
The Bland-Altman plot comparing the ground truth to the computed outcome is depicted in Fig. 7(a). In this case, the limits of agreement (LOA) ranged from $[-4.04,3.80]$ beats/min, with 95% of the differences falling within this interval. Fig. 7(b) shows the scatter plot comparing the ground-truth HR values to those calculated for 16 test subjects (12 undergoing cycling and 4 on treadmill exercises). Additionally, the line of best fit, represented as $y=1.005x-0.7492$, is displayed, where $x$ is the ground-truth HR and $y$ is the calculated HR. The Pearson correlation coefficient $R$ achieved using the proposed AM-GAN is 0.9970.
@@
Figure 9: RR results, measured in breaths/minute (bpm), were derived from the datasets of 16 test subjects during cycling and treadmill exercises. Displayed are (a) a Bland-Altman plot and (b) a scatter plot illustrating the Pearson correlation.
The average absolute error ($Err1$), the average absolute error percentage ($Err2$) and the standard deviation ($SD$) are defined to further evaluate the accuracy of the proposed AM-GAN by performing HR and RR calculations. Table 1 presents the $Err1$ and $Err2$ values for the cycling and treadmill exercise. The average across the 16 testing subject datasets, $\overline{Err1}$ of the proposed AM-GAN is $1.37\pm 0.32$ beats/min (Mean $\pm$ SD), and that of $\overline{Err2}$ is $1.29\%\pm 0.35\%$ for the HR calculation. $\overline{Err1}$ is $2.49\pm 0.52$ breaths/min (Mean $\pm$ SD) and $\overline{Err2}$ is $10.78\%\pm 1.94\%$ for the RR calculation.
@@
Table 1: Average Absolute Error ($Err1$) (in beats/min) and Average Error Percentage ($Err2$) (%) for all subjects (F-female, M-male, T-treadmill, C-cycling) of the in-house LU dataset during the cycling and treadmill exercises.
HR calculation on IEEE-SPC dataset: The performance of AM-GAN was assessed using the IEEE SPC dataset as a benchmark, which includes 12 training datasets and 10 testing datasets, detailed in Table 2. $Err1$, $Err2$, and $SD$ were calculated and compared with results from other methods using a testing dataset of 10 subjects. This dataset included data collected under strenuous exercise conditions such as weight-lifting, swimming, boxing, and handshaking. In these scenarios, the prominence of MAs can make the assessment of physiological parameters more difficult. This table shows that, for 10 test subjects with the most noise, the AM-GAN produced HR calculations with $Err1$ of 1.81 and $SD$ of 0.76 beats/min, and also with $Err2$ of 1.86 and $SD$ of 1.63 %. From the $Err1$ and $Err2$ values, the AM-GAN has a better performance compared to the other traditional de-noised methods, but it is not yet as accurate as the CorNET [7] network that adopts a deep learning architecture. CorNET and AM-GAN use different forms of output; the input of CorNET is the PPG signal and the output is the HR value, which is more of a black-box structure and does not clearly reflect the characteristics of the signals corresponding to different HRs. On the contrary, the AM-GAN has an input of PPG signal with noise and an output of noise-free PPG signal, which is used to calculate HR as well as other physiological parameters using the noise-free PPG signal. Additionally, the difference in $Err1$ between the proposed AM-GAN and the reference method (MR) suggests a fundamental similarity in their performance. Meanwhile, the $SD$ of the proposed method ($SD=0.76$ beats/min) is smaller than that of the MR method ($SD=1.42$ beats/min), it signifies a reduced intergroup error value for the proposed AM-GAN, indicative of superior robustness. Furthermore, the proposed AM-GAN incorporates an Attention Mechanism ($Err1=1.81$ beats/min), enhancing the precision of physiological parameter calculation, in contrast to our previous PPG-GAN method without utilising such a mechanism ($Err1=1.93$ beats/min).
Subject Activity TROIKA [4] JOSS [18] SPECMAR [19] MURAD [20] Temko [21] ANFA [22] MR method [15] CorNET[7] PPG-GAN [23] Proposed AM-GAN
1 T2 6.63(8.76) 8.07(10.90) 6.57 6.35(7.99) 9.59(10.90) 4.99(7.52) 5.69(7.63) 1.60 3.67 3.46(6.29)
2 T2 1.94(2.56) 1.61(2.01) 1.76 1.15(1.58) 3.65(2.43) 1.46(1.99) 1.02(1.35) 0.24 1.35 1.21(1.45)
3 T3 1.35(1.04) 3.10(2.69) 2.28 1.51(1.25) 3.90(1.51) 1.77(0.94) 1.02(0.83) 1.60 1.53 1.46(0.89)
4 T3 7.82(4.88) 7.01(4.49) 2.77 2.96(1.87) 2.44(1.49) 2.89(1.51) 1.89(1.27) 2.04 2.75 2.66(1.53)
5 T3 2.46(2.00) 2.99(2.52) 2.94 2.78(2.24) 2.14(1.70) 2.93(0.93) 1.83(1.55) 0.95 2.25 2.09(1.81)
6 T3 1.73(1.27) 1.67(1.23) 4.80 1.54(1.12) 2.60(0.90) 1.98(1.01) 1.15(0.85) 0.28 1.44 1.31(0.99)
7 T2 3.33(3.90) 2.80(3.46) 2.72 2.03(2.24) 1.86(1.78) 2.01(3.65) 1.82(2.10) 0.28 1.96 1.73(2.45)
8 T3 3.41(2.43) 1.88(1.32) 3.28 2.02(1.61) 0.85(1.96) 1.92(1.31) 1.73(1.22) 0.67 1.77 1.68(1.21)
9 T3 2.69(2.12) 0.92(0.74) 1.55 0.97(0.76) 3.06(0.80) 1.86(1.04) 0.99(0.80) 0.42 1.58 1.52(1.13)
10 T2 0.51(0.59) 0.49(0.57) 0.82 0.83(0.66) 3.38(0.59) 1.16(0.97) 0.84(0.98) 0.57 0.99 0.93(0.90)
Mean 1-10 3.19(2.96) 3.05(2.99) 2.95 2.21(2.13) 3.35(2.41) 2.30(2.09) 1.80(1.85) 0.86 1.93 1.81(1.86)
SD 1-10 2.32(2.41) 2.53(3.04) 1.67 1.62(2.13) 2.37(3.04) 1.99(2.08) 1.42(2.06) 0.65 0.79 0.76(1.63)
@@
Cross-dataset HR validation: Generalisation capacity is crucial for real-time physiological monitoring. To validate this aspect, cross-database evaluations have been performed to assess the generalisation ability of the proposed AM-GAN. The AM-GAN is trained on the in-house LU dataset and tested on the PPGDalia dataset for HR calculation. The proposed AM-GAN is compared with our previous adversarial network model without attention mechanism, i.e., PPG-GAN [23]. The cross-dataset HR calculation results of the proposed AM-GAN and the PPG-GAN are demonstrated in Table 3.
S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 S11 S12 S13 S14 S15 Mean (SD)
DeepPPG [17] 7.73 6.74 4.03 5.90 18.51 12.88 3.91 10.87 8.79 4.03 9.22 9.35 4.29 4.37 4.17 7.65 (4.15)
NAS-PPG [24] 5.46 5.01 3.74 6.48 12.68 10.52 3.31 8.07 7.91 3.29 7.05 6.76 3.84 4.85 3.57 6.17 (2.77)
Q-PPG [25] 4.29 3.62 2.44 5.73 10.33 5.26 2.00 7.09 8.60 3.09 4.99 6.25 1.92 3.02 3.55 4.81 (2.46)
Aug-TEMPONet [26] 4.97 4.34 2.39 6.14 9.41 3.63 2.23 9.14 10.98 3.40 5.27 7.64 2.05 2.84 3.61 5.20 (2.86)
Ours
PPG-GAN [23] 4.81 3.85 2.62 5.31 9.63 4.69 2.32 9.99 8.06 3.54 3.93 6.68 1.84 3.42 2.95 4.91 (2.57)
AM-GAN 3.01 2.98 2.04 3.97 7.9 4.64 2.41 7.12 5.83 2.33 3.09 6.38 1.73 2.49 2.01 3.86 (2.02)
@@
Table 3: $Err1$ values for cross-dataset HR calculation (training on LU dataset and testing on PPGDalia dataset) by the proposed AM-GAN and PPG-GAN method to compare with the SOTA methods.
@@
The comprehensive cross-dataset HR calculation outcomes for both the proposed AM-GAN and the PPG-GAN are presented in Table 3. The proposed AM-GAN illustrated the best results, given mean HR calculations with $\overline{Err1}$ of $3.86$ and $SD$ of $2.02$ beats/min, compared with PPG-GAN and SOTA methods. However, subjects S5, S6 and S8 have posed challenges with high $Err1$ values. Comparison of S5, S6 and S8 shows that the underlying causes of high errors are different. S5 and S6 are primarily affected by the out-of-distribution samples, and S8 is mainly affected by the lower signal quality of collected data. In summary, the cross-dataset HR validations indicate that the proposed AM-GAN exhibits robust generalisation capabilities in scenarios characterised by unknown noise and MAs.

### v2

@@ 2.2.2 Cross-attention Mechanism for Multi-sensor Fusion
A self-attention mechanism was proposed as a core component of the transformer model [b30ashish2017attention]. The self-attention accepts a tensor comprising sequential data and establishes its self-correlation. Specifically, it involves linear projections of the input $X$ into queries ($Q$), keys ($K$), and values ($V$). The scaled dot-product attention ($A$) is then applied to these projections as $A=softmax((QK^{T})/\sqrt{d_{k}})V$, here $d{{}_{k}}$ is the dimensionality of the queries and keys.
In contrast to the self-attention mechanism that correlates the input $X$ with itself, the multi-head cross-attention mechanism is proposed for multi-sensor fusion. They are two inputs $X_{1}$ and $X_{2}$, where $X_{1}$ represents query $Q$ tensors and $X_{2}$ represents the key $K$ and value $V$ tensors. In the proposed attention mechanism, these tensors are related to different sensor inputs, including encoded noisy PPG and encoded triaxial ACC with related VEL feature maps after four downsample blocks of the generator. The encoded feature maps are fed into the attention mechanism, as shown in Fig. 2 and Algorithm 1. Algorithm 1 presents the pseudocode implementation of our proposed multi‑sensor fusion cross‑attention mechanism. Specifically, the proposed attention mechanism takes the generated PPG encoded feature map processed by a (feature pooling) $1\times 1$ convolution layer to create the query $Q$. Also, the key $K$ and value $V$ pair is formed through a similar $1\times 1$ convolution layer of generated combined encoded feature map with triaxial ACC and the related VEL. The $K$ and $V$ vectors are used to calculate the attention weights, indicating to the model which parts of the MAs to focus on. The $Q$ vector is the information that these attention weights will be applied to, with the aim of highlighting the MA features in the PPG encoded feature map that are most relevant to the ACC and VEL combined encoded feature map.
@@
Algorithm 1 Multi-sensor fusion cross-attention mechanism pseudocode
1:#\# encoder network $g_{enc}$, decoder network $g_{dec}$
2:#\# query $Q$, key $K$ and value $V$
[…]
4:for each minibatch $x_{PPG},x_{ACCVEL}$ in loader do
5:  $x_{PPG},x_{ACCVEL}\leftarrow\mathsf{preprocess}(x_{PPG}),\;\mathsf{preprocess}(x_{ACCVEL})$
6:  $X_{PPG},X_{ACCVEL}\leftarrow g_{enc}(x_{PPG}),\,g_{enc}(x_{ACCVEL})$ #\# encoder outputs
7:  $\leftarrow conv_{1\times 1}(X_{PPG}) #\#V,KV,K$ outputs
8:  $\leftarrow conv_{1\times 1}(X_{ACCVEL}) #\#QQ$ outputs
9:  $W_{MA}\leftarrow softmax((QK^{T})/\sqrt{d_{k}})$ #\# MA attention weights
10:  $S\leftarrow X_{PPG}-W_{MA}V$ #\# MA-removal signal features
11:  $s_{PPG}\leftarrow g_{dec}(S)$ #\# MA-removal PPG signal
@@ 2.4 Loss function
The discriminator binary cross entropy loss function is [b20ronneberger2015u]
@@ 3 Selective Measurement Protocols
Four datasets provided by Loughborough University (in-house LU dataset), the IEEE Signal Processing Cup 2015 (IEEE-SPC dataset) and the publicly available PPG dataset (PPGDalia dataset) with their ethical approvals are experimented with the proposed AM-GAN.
@@
The in-house multi-wavelength C2 datasets were collected using a protocol implemented in a controlled environmental chamber, where measurements (at rest and in motion) were taken at specific simulated altitudes at three different Oxygen levels of 21%, 18% and 16%. During cycling exercise, the subjects performed the exercise at three different intensities; each stage lasted four minutes with a recovery of 60 seconds in between. The data was simultaneously gathered from the multi-wavelength C2 devices at a sampling frequency of 1024 $Hz$ worn on the wrist and attached on the chest, from four subjects aged 25 - 35 years and with type I, type II, type IV, and type VI skin tones (Fitzpatrick scale). In the cycling exercise, continuous SpO2 reference data were obtained using a Finapres® NOVA patient monitor.
The IEEE-SPC dataset features a two-channel PPG signal using green (515 nm) LEDs, a triaxial ACC signal, and an ECG reference, collected from 12 male subjects aged between 18 and 35 years. These datasets were sampled at a frequency of 125 Hz [b6zhang2014troika]. The IEEE-SPC includes the training and testing datasets. In the training datasets, 12 subjects ran at varying speeds. Furthermore, 10 subjects (aged 19-58 years) performed intensive arm movements, i.e., boxing, in the testing datasets.
The PPGDalia includes synchronised recordings of ECG, wrist-worn PPG, and acceleration data from 15 subjects. Each subject contributed approximately two hours of recorded data. A consistent time shift between PPG and triaxial ACC signals was identified and manually fixed [b29reiss2019deep].
@@ 4.1 Experimental Setup
An eight-second sliding window technique is applied to analyse both the input and reference PPG signals, advancing one second at each step. All input and reference PPG signals are downsampled to 32 $Hz$. During the experiments, the proposed AM-GAN undergoes training for 100 epochs, with both the generator and discriminator networks being optimised concurrently using the Adam optimiser at a learning rate of 0.0002. The input PPG data is pre-processed with an eighth-order Butterworth band-pass filter with band edges set at 0.2 $Hz$ and 6.5 $Hz$. We apply one-sided label smoothing to the discriminator’s real targets: during training, the ground-truth label for real images is set to 0.9 instead of 1.0. This small perturbation prevents the discriminator from becoming over-confident and keeps its gradients in a meaningful range. In all experiments, HR and RR are computed using the extraction algorithms defined in [b21zheng2023rapid].
@@ 4.2 Experimental Results
Waveform quality on in-house LU dataset: Fig. 4 shows the selected PPG waveform in the time domain from the different physical activity intensities (subject-M10, treadmill), i.e., rest, low-intensity (6km/hkm/h), medium-intensity (9km/hkm/h), and high-intensity (12km/hkm/h). The generated PPG signals illustrate that the AM-GAN is capable of yielding results that closely approximate the reference signals across various levels of exercise intensity, thereby achieving satisfactory outcomes.
Figure 4: A representative example for the waveform comparison of the original noisy signal, the reference signal and the generated signal at different physical activity intensities; i.e., (a) at rest, (b) low-intensity activity (6km/hkm/h), (c) medium-intensity activity (9km/hkm/h), and (d) high-intensity activity (12km/hkm/h).
$R=\frac{\int_{0}^{T}G(p(t),a_{x,y,z}(t))s_{ref}(t)dt}{\sqrt{\int_{0}^{T}G(p(t),a_{x,y,z}(t))^{2}dt}\sqrt{\int_{0}^{T}s_{ref}(t)^{2}dt}}$ (7)
To validate the error of the PPG signal generated by the AM-GAN method with respect to the reference PPG signal, the Pearson correlation ($R$) was applied for quantifying waveform, as shown in (7), where $G(p(t),a_{x,y,z}(t))$ is the zero-mean generated PPG signal and $s_{ref}(t)=MR(p(t),a_{x,y,z}(t))$ is the zero-mean reference signal, for each vector $p(t)$ in the test set, $\mathcal{P}_{Te}$. An eight-second sliding time window was employed for the calculation of the waveform quality. The waveform quality index ($R$), for the treadmill exercise test cases of subject F01, is shown in Fig. 5(a). The mean Pearson correlation ($\overline{\langle R\rangle}$) for all LU testing subjects is illustrated in Fig. 5(b). It is clear that the AM-GAN outputs are highly correlated with the reference signals with a high level of consistency across the entire set of test vectors and for all testing subjects, and the $Mean$ value of all testing subjects’ waveform quality ($\overline{\langle R\rangle}$) is 0.9522.
@@
The outcomes of continuous HR monitoring during treadmill and cycling exercises are shown in Fig. 6(a) and 6(b), where they are compared to the reference values from a commercial device. In Fig. 6, Polar refers to the reference heart rate recorded with a Polar Bluetooth Smart chest strap. In both figures, the MA-removal reference uses our previously published algorithm [b21zheng2023rapid]. The accuracy in capturing HR is consistent and closely matches the provided reference.
@@
The Bland-Altman plot comparing the ground truth to the computed outcome is depicted in Fig. 7(a). In this case, the limits of agreement (LOA) ranged from [4.04,3.80][-4.04,3.80] beats/min, with 95% of the differences falling within this interval. Fig. 7(b) shows the scatter plot comparing the ground-truth HR values to those calculated for 16 test subjects (12 undergoing cycling and 4 on treadmill exercises). Additionally, the line of best fit, represented as y=1.005x0.7492y=1.005x-0.7492, is displayed, where $x$ is the ground-truth HR and $y$ is the calculated HR. The Pearson correlation coefficient $R$ achieved using the proposed AM-GAN is 0.9970.
@@
Figure 9: $R$ results, measured in breaths/minute (bpm), were derived from the datasets of 16 test subjects during cycling and treadmill exercises. Displayed are (a) a Bland-Altman plot and (b) a scatter plot illustrating the Pearson correlation.
The mean absolute error ($Err1$), the mean absolute error percentage ($Err2$) and the standard deviation ($SD$) are defined to further evaluate the accuracy of the proposed AM-GAN by performing HR and RR calculations. Table 1 presents the $Err1$ and $Err2$ values for the cycling and treadmill exercise. The average across the 16 testing subject datasets, $\overline{Err1}$ of the proposed AM-GAN is $1.37\pm 0.32$ beats/min (Mean $\pm$ SD), and that of $\overline{Err2}$ is $1.29\%\pm 0.35\%$ for the HR calculation. $\overline{Err1}$ is $2.49\pm 0.52$ breaths/min (Mean $\pm$ SD) and $\overline{Err2}$ is $10.78\%\pm 1.94\%$ for the RR calculation.
Table 1: Mean Absolute Error ($Err1$) (in beats/min) and Average Error Percentage ($Err2$) (%) for all subjects (F-female, M-male, T-treadmill, C-cycling) of the in-house LU dataset during the cycling and treadmill exercises.
@@
HR calculation on IEEE-SPC dataset: The performance of AM-GAN was assessed using the IEEE SPC dataset as a benchmark, which includes 12 training datasets and 10 testing datasets, detailed in Table 2. $Err1$, $Err2$, and $SD$ were calculated and compared with results from other methods using a testing dataset of 10 subjects. This dataset included data collected under strenuous exercise conditions such as weight-lifting, swimming, boxing, and handshaking. In these scenarios, the prominence of MAs can make the assessment of physiological parameters more difficult. This table shows that, for 10 test subjects with the most noise, the AM-GAN produced HR calculations with $Err1$ of 1.81 and $SD$ of 0.76 beats/min, and also with $Err2$ of 1.86 and $SD$ of 1.63 %. From the $Err1$ and $Err2$ values, the AM-GAN has a better performance compared to the other traditional de-noised methods, but it is not yet as accurate as the CorNET [b9biswas2019cornet] network that adopts a deep learning architecture. CorNET and AM-GAN use different forms of output; the input of CorNET is the PPG signal and the output is the HR value, which is more of a black-box structure and does not clearly reflect the characteristics of the signals corresponding to different HRs. On the contrary, the AM-GAN has an input of PPG signal with noise and an output of noise-free PPG signal, which is used to calculate HR as well as other physiological parameters using the noise-free PPG signal. Additionally, the difference in $Err1$ between the proposed AM-GAN and the reference method (MR) suggests a fundamental similarity in their performance. Meanwhile, the $SD$ of the proposed method ($SD=0.76$ beats/min) is smaller than that of the MR method ($SD=1.42$ beats/min), it signifies a reduced intergroup error value for the proposed AM-GAN, indicative of superior robustness. Furthermore, the proposed AM-GAN incorporates an Attention Mechanism ($Err1=1.81$ beats/min), enhancing the precision of physiological parameter calculation, in contrast to our previous PPG-GAN method without utilising such a mechanism ($Err1=1.93$ beats/min).
@@
Subject Activity TROIKA [b6zhang2014troika] JOSS [b25zhang2015photoplethysmography] SPECMAR [b26islam2019specmar] MURAD [b27chowdhury2016real] Temko [b28temko2017accurate] ANFA [b18zheng2022adaptive] MR method [b21zheng2023rapid] CorNET [b9biswas2019cornet] PPGANET [b33sawangjai2025removal] MAICR [b34hao2024ppg] EPDA [b36bradley2024opening] PPG-GAN [b29zheng2022ppg] Proposed AM-GAN
1 T2 6.63(8.76) 8.07(10.90) 6.57 6.35(7.99) 9.59(10.90) 4.99(7.52) 5.69(7.63) 1.60 3.79 4.36 6.41 3.67 3.46(6.29)
2 T2 1.94(2.56) 1.61(2.01) 1.76 1.15(1.58) 3.65(2.43) 1.46(1.99) 1.02(1.35) 0.24 1.47 1.06 1.82 1.35 1.21(1.45)
3 T3 1.35(1.04) 3.10(2.69) 2.28 1.51(1.25) 3.90(1.51) 1.77(0.94) 1.02(0.83) 1.60 1.58 2.57 2.68 1.53 1.46(0.89)
4 T3 7.82(4.88) 7.01(4.49) 2.77 2.96(1.87) 2.44(1.49) 2.89(1.51) 1.89(1.27) 2.04 2.83 1.32 3.01 2.75 2.66(1.53)
5 T3 2.46(2.00) 2.99(2.52) 2.94 2.78(2.24) 2.14(1.70) 2.93(0.93) 1.83(1.55) 0.95 2.42 1.89 2.52 2.25 2.09(1.81)
6 T3 1.73(1.27) 1.67(1.23) 4.80 1.54(1.12) 2.60(0.90) 1.98(1.01) 1.15(0.85) 0.28 1.69 1.66 1.79 1.44 1.31(0.99)
7 T2 3.33(3.90) 2.80(3.46) 2.72 2.03(2.24) 1.86(1.78) 2.01(3.65) 1.82(2.10) 0.28 1.99 2.41 1.98 1.96 1.73(2.45)
8 T3 3.41(2.43) 1.88(1.32) 3.28 2.02(1.61) 0.85(1.96) 1.92(1.31) 1.73(1.22) 0.67 1.81 1.84 2.01 1.77 1.68(1.21)
9 T3 2.69(2.12) 0.92(0.74) 1.55 0.97(0.76) 3.06(0.80) 1.86(1.04) 0.99(0.80) 0.42 1.86 2.63 2.22 1.58 1.52(1.13)
10 T2 0.51(0.59) 0.49(0.57) 0.82 0.83(0.66) 3.38(0.59) 1.16(0.97) 0.84(0.98) 0.57 1.03 0.81 1.36 0.99 0.93(0.90)
Mean 1-10 3.19(2.96) 3.05(2.99) 2.95 2.21(2.13) 3.35(2.41) 2.30(2.09) 1.80(1.85) 0.86 2.04 2.05 2.58 1.93 1.81(1.86)
SD 1-10 2.32(2.41) 2.53(3.04) 1.67 1.62(2.13) 2.37(3.04) 1.99(2.08) 1.42(2.06) 0.65 0.79 1.01 1.43 0.79 0.76(1.63)
@@
Cross-dataset HR validation: Generalisation capacity is crucial for real-time physiological monitoring. To validate this aspect, cross-database evaluations have been performed to assess the generalisation ability of the proposed AM-GAN. The AM-GAN is trained on the in-house LU dataset and tested on the PPGDalia dataset for HR calculation. The proposed AM-GAN is compared with our previous adversarial network model without an attention mechanism, i.e., PPG-GAN [b29zheng2022ppg]. The cross-dataset HR calculation results of the proposed AM-GAN and the PPG-GAN are demonstrated in Table 3.
@@
Table 3: $Err1$ values for cross-dataset HR calculation (training on LU dataset and testing on PPGDalia dataset) by the proposed AM-GAN and PPG-GAN method to compare with the SOTA methods. p-values for a two-sided Wilcoxon signed-rank test against the AM-GAN model.
S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 S11 S12 S13 S14 S15 Mean (SD) p-value
DeepPPG [b29reiss2019deep] 7.73 6.74 4.03 5.90 18.51 12.88 3.91 10.87 8.79 4.03 9.22 9.35 4.29 4.37 4.17 7.65 (4.15) 6.08e-5
NAS-PPG [b30risso2021robust] 5.46 5.01 3.74 6.48 12.68 10.52 3.31 8.07 7.91 3.29 7.05 6.76 3.84 4.85 3.57 6.17 (2.77) 6.10e-5
Q-PPG [b31burrello2021q] 4.29 3.62 2.44 5.73 10.33 5.26 2.00 7.09 8.60 3.09 4.99 6.25 1.92 3.02 3.55 4.81 (2.46) 1.53e-3
Aug-TEMPONet [b32burrello2022improving] 4.97 4.34 2.39 6.14 9.41 3.63 2.23 9.14 10.98 3.40 5.27 7.64 2.05 2.84 3.61 5.20 (2.86) 8.54e-4
PPGANET [b33sawangjai2025removal] 4.51 3.37 2.33 4.25 8.68 4.76 2.37 7.04 7.75 2.80 5.19 6.08 2.09 3.02 3.19 4.49 (2.08) 2.01e-3
Adaptive Q-PPG [b38kechris2024kid] 3.80 3.50 2.07 5.18 5.76 3.38 1.59 6.99 8.94 2.75 3.37 5.61 1.49 2.45 2.80 3.98 (2.12) 7.76e-1
PULSE [b37kasnesis2023feature] 4.75 3.31 2.22 5.25 7.43 4.22 2.28 8.93 6.95 2.93 3.98 6.57 1.70 3.22 2.88 4.44 (2.16) 1.03e-2
PPG-GAN [b29zheng2022ppg] 4.81 3.85 2.62 5.31 9.63 4.69 2.32 9.99 8.06 3.54 3.93 6.68 1.84 3.42 2.95 4.91 (2.57) 1.83e-4
AM-GAN 3.01 2.98 2.04 3.97 7.90 4.64 2.41 7.12 5.83 2.33 3.09 6.38 1.73 2.49 2.01 3.86 (2.02)
@@
The comprehensive cross-dataset HR calculation outcomes for both the proposed AM-GAN and the PPG-GAN are presented in Table 3. The proposed AM-GAN illustrated the best results, given mean HR calculations with $\overline{Err1}$ of $3.86$ and $SD$ of $2.02$ beats/min, compared with PPG-GAN and SOTA methods. However, subjects S5, S6 and S8 have posed challenges with high $Err1$ values. Comparison of S5, S6 and S8 shows that the underlying causes of high errors are different. S5 and S6 are primarily affected by the out-of-distribution samples, and S8 is mainly affected by the lower signal quality of collected data. Pairwise comparisons using the two-sided Wilcoxon signed-rank test with Bonferroni correction indicated that the proposed AM-GAN method significantly outperforms DeepPPG, NAS-PPG, Q-PPG, Aug-TEMPONet, PPGANET, and PPG-GAN ($p<0.05$). No statistically significant differences were found between AM-GAN and Adaptive Q-PPG or PULSE. In summary, the cross-dataset HR validations indicate that the proposed AM-GAN exhibits robust generalisation capabilities in scenarios characterised by unknown noise and MAs.
Table 4: Comparison of $Err1(Err2)$ (in % SpO2 and % error) for cross-dataset SpO2 estimation. Training was done on LU dataset and testing on C2 dataset under cycling (C) condition.
Subject S1 S1 S1 S2 S2 S2 S3 S3 S3 S4 S4 S4 Mean $\pm$ SD
Oxygen Level 21% 18% 16% 21% 18% 16% 21% 18% 16% 21% 18% 16%
[…]
Cross-dataset SpO2 validation: For an in-depth exploration of the AM-GAN’s robustness, cross-dataset SpO2 validation results are outlined in Table 4. Notably, given that the in-house C2 dataset incorporates a multi-wavelength opto-physiological monitoring sensor, the cross-dataset SpO2 validation necessitates the utilisation of two additional wavelength channels, i.e., red and infrared channels. The relationship between SpO2 and the ratio of absorbance $\rho$ [b21zheng2023rapid] is provided by the C2 experiments.
$\textrm{SpO}_{2}\%=109-21.5\rho$ (8)
where $\rho=(AC_{Red}/DC_{Red})/(AC_{IR}/DC_{IR})$.
From Table 4, the AM-GAN performance ($\overline{Err1}(\overline{Err2})$ of $1.65(1.79)$ and $SD$ of 0.79(0.99)0.79(0.99)%) consistently outperforms our prior approach, which lacked an attention mechanism, across all evaluation metrics. Fig. 10 demonstrated that the calculated SpO2 from AM-GAN closely aligns with the reference SpO2 in three different oxygen levels. The performance of SpO2 calculation exhibits commendable alignment with the reference SpO2, revealing the $Err1$ of 1.651.65% for cycling exercises in three different oxygen levels. The cross-dataset SpO2 validations indicate that the proposed framework exhibits robust generalisation capabilities in scenarios characterised by unknown noise and motion artefacts.
Figure 10: SpO2 calculation results from C2 subject 4 (three oxygen levels) in cycling exercise. The X-axis illustrates the cycling activity intensities.